# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ 5.5 TFB Beyond the Naive LoRA
Our proposed TFB is compatible with various LoRA-based methods that use different initialization strategies (Meng et al., 2024), parameter sharing schemes (Kopiczko et al., 2023), and optimization approaches (Zhang et al., 2023). We evaluate TFB on two representative LoRA variants:
  • VeRA (Kopiczko et al., 2023): Uses shared low-rank matrices ${\bm{B}}$ and ${\bm{A}}$ across layers, with layer-specific trainable scalar vector ${\bm{d}}$ and bias vector ${\bm{b}}$.
  • PiSSA (Meng et al., 2024): Employs an alternative initialization while maintaining LoRA’s training process.
We conduct experiments on Llama-3.1-8B following the protocol described in Sec. 5.4. Detailed results for individual datasets and the technical adaptations required to satisfy the assumptions in Eqn. 3 (variational distribution to be a low-rank isotropic Gaussian) for each LoRA variant are provided in Appendix C.4.
The results in Table 4 demonstrate TFB’s broad applicability: it substantially reduces ECE across all LoRA-like PEFT methods (e.g., from 5.11 to 1.44 for VeRA). Importantly, it maintains strong ACC and NLL scores with minimal performance degradation, validating the effectiveness of low-rank isotropic Gaussian distribtuions for variational inference of LLMs.
Table 4: Performance of different LoRA-like PEFT methods on the combined dataset of six commonsense reasoning tasks. Boldface denotes the best performance.
@@
Table 5: Performance of Last-Layer TFB (LL TFB) applied to the combined dataset of 6 commonsense reasoning tasks on Llama3.1-8B pre-trained weights. Boldface denotes the best performance.
@@
5.6 Improving the Efficiency of TFB
Inspired by Harrison et al. (2024), we adopt last-layer Bayesianization to increase the inference speed. In our approach, we employ the same prior standard deviation $\sigma_{q}$ as identified through full-model Bayesian tuning, but restrict the Bayesian treatment to the last layer of model. By limiting inference to sampling from the posterior distribution of the last layer, the outputs of earlier layers can be reused, significantly improving inference speed. Although restricting Bayesianization to the last layer sacrifices some ability to effectively model uncertainty across the entire model, this approach allows for a higher number of posterior samples, resulting in more accurate approximations.
As shown in Table 5, when only $N=10$ samples are used, the uncertainty estimation of last-layer Bayesianization performs worse than that of full-model Bayesianization. Nevertheless, the improved posterior estimation with $N=100$ samples enables last-layer Bayesianization to achieve better performance. For a comprehensive analysis of its performance, please refer to Appendix C.5.
@@ 6 Conclusion
In this paper, we introduce Training-Free Bayesianization (TFB), a novel framework that transforms trained LoRA adapters into Bayesian ones without additional training. By systematically searching for the maximally acceptable variance in the weight posterior within a family of low-rank isotropic Gaussian distributions, TFB provides a practical solution to uncertainty estimation in LLMs. We theoretically demonstrate that TFB’s variance maximization process is equivalent to variational inference under mild conditions, while empirically showing its superior performance across various settings and model configurations. Our framework’s simplicity and effectiveness, requiring only an anchor dataset for search, makes it widely applicable across different domains. As LLMs continue to evolve, TFB represents a significant step toward more reliable and uncertainty-aware AI systems, paving the way for future research in adaptive and trustworthy machine learning.
7 Limitations
TFB is subject to several limitations. First, our approach relies on the availability of an anchor dataset for determining search range and stopping criteria. Although this dataset doesn’t require supervision or prior use in LoRA training, its quality and representativeness could impact the effectiveness of uncertainty estimation. Second, by constraining the family of full-weight posteriors to low-rank isotropic Gaussian distributions, TFB may not capture more complex uncertainty patterns that could be present in the data. While this constraint enables our training-free approach, it represents a trade-off between computational efficiency and model expressiveness. Finally, while we have demonstrated the effectiveness of TFB in various settings, its performance in more complex generation tasks requires further investigation. Future work could explore extending the framework to handle more sophisticated language generation scenarios and broader applications.
@@ 7 Limitations
In Appendix A, we provide detailed proofs for all theorems presented in the main paper. In Appendix B, we describe our experimental methodology. Finally, in Appendix C, we present additional empirical results, including the full results of experiments with different searching strategies for optimal $\sigma_{q}^{*}$, variational distribution family, LLM backbones, and LoRA-like PEFTs.
Appendix A Proof of Theorems
Theorem 4.1.
With the pre-trained weight matrix ${\bm{W}}_{0}\in\mathbb{R}^{m\times n}$, the low-rank weight update matrix $\{{\bm{B}}^{\prime}\in\mathbb{R}^{m\times r},{\bm{A}}^{\prime}\in\mathbb{R}^{r\times n}\}$ transformed from the given matrices $\{{\bm{B}},{\bm{A}}\}$ following Eqn. 4 and 5, suppose that the variational distribution of ${\bm{A}}^{\prime}$ is Gaussian $q({\bm{A}}^{\prime}|{\bm{\theta}})=\prod_{ij}{\mathcal{N}}(A_{ij}|M_{ij},\Omega_{ij}^{2})$, where ${\bm{M}}=[M_{ij}=A^{\prime}_{ij}]\in\mathbb{R}^{r\times n}$ is its mean and ${\bm{\Omega}}=[\Omega_{ij}]\in\mathbb{R}^{r\times n}$ is the standard deviation calculated as in Eqn. 6. The equivalent variational distribution $q(\operatorname{vec}({\bm{W}})|\sigma_{q})$ defined on the full weight matrix ${\bm{W}}$ is
${\bm{\Sigma}}_{q}$ $=\sigma_{q}^{2}\cdot{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &{\bm{O}}_{m-r}\end{bmatrix}.$
Proof.
We have the following lemma from BLoB that calculates the covariance matrix of a given low-rank Bayesianization scheme $\{{\bm{B}},{\bm{A}},{\bm{\Omega}}\}$ (Wang et al., 2024d).
Lemma A.1.
@@ Appendix A Proof of Theorems
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}]\cdot[\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\top}].$
Based on the assumption outlined in Eqn. 4, 5, and 6, we have the following properties about ${\bm{B}}^{\prime}$, ${\bm{M}}$, and ${\bm{\Omega}}$ of TFB:
@@
where ${\bm{U}}^{\top}{\bm{U}}={\bm{I}}_{r},{\bm{U}}{\bm{U}}^{\top}=\begin{bmatrix}{\bm{I}}_{r}&\\ &{\bm{O}}_{m-r}\end{bmatrix}.$ (17)
@@
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}^{\prime}]\cdot[\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\prime\top}]$ (18)
$=[{\bm{I}}_{n}\otimes{\bm{B}}^{\prime}]\cdot[{\bm{I}}_{n}\otimes\operatorname{diag}(1/{\bm{d}})^{2}]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\prime\top}]$ (19)
@@
$=\sigma_{q}^{2}\cdot{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &{\bm{O}}_{m-r}\end{bmatrix},$ (22)
@@
Theorem 4.2.
Suppose the evaluation metric $l_{\mathcal{D}}(\sigma_{q})$ defined following Assumption 4.1 is locally convex within the range of $0<\sigma_{q}<\epsilon_{0}$. Suppose the posterior distribution of ${\bm{W}}$ given $\sigma_{q}$ is defined following Theorem 4.2. Suppose we have the prior distribution $P(\operatorname{vec}({\bm{W}}))={\mathcal{N}}(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{p},{\bm{\Sigma}}_{p})$, where ${\bm{\mu}}_{p}={\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}})$, and ${\bm{\Sigma}}_{p}=\sigma_{p}^{2}{\bm{I}}$ with $\sigma_{p}\gg\epsilon_{0}$. Then $\exists\epsilon>0$, s.t. the two following optimization problems
(i) Minimization of Variational Free-Energy (Blundell et al., 2015)
$\min_{\sigma_{q}}\quad l_{\mathcal{D}}(\sigma_{q})+\lambda\cdot\operatorname{KL}[q({\bm{W}}|\sigma_{q})\parallel P({\bm{W}})],$ (23)
@@
s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\epsilon,$
are equivalent, where $\lambda$ is the coefficient of the KL-divergence regularization.
@@
First we prove the KL divergence term is convex w.r.t. $\sigma_{q}$. For two Gaussian distributions $q$ and $p$ whose covariance matrices ${\bm{\Sigma}}_{q}\in\mathbb{R}^{d\times d}$ and ${\bm{\Sigma}}_{p}\in\mathbb{R}^{d\times d}$ are both full-rank, and their means as ${\bm{\mu}}_{q}\in\mathbb{R}^{d}$ and ${\bm{\mu}}_{p}\in\mathbb{R}^{d}$, we have their KL-divergence as
@@
$q(\operatorname{vec}({\bm{W}})|\sigma_{q})$ $={\mathcal{N}}\left(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}}),{\bm{\Sigma}}_{q}=\sigma_{q}^{2}\cdot{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &{\bm{O}}_{m-r}\end{bmatrix}\right),$ (26)
$P(\operatorname{vec}({\bm{W}})|\sigma_{p})$ $={\mathcal{N}}\left(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{p}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}}),{\bm{\Sigma}}_{p}=\sigma_{p}^{2}\cdot{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &{\bm{O}}_{m-r}\end{bmatrix}\right).$
Substituting Eqn. 26 back into Eqn. 25, we have
$\operatorname{KL}[q(\operatorname{vec}({\bm{W}})|\sigma_{q})\|P(\operatorname{vec}({\bm{W}}|\sigma_{p}))]$ $=\tfrac{nr}{2}\cdot\left[\log(\sigma_{p}^{2})-1+\left\{-\log(\sigma_{q}^{2})+\tfrac{\sigma_{q}^{2}}{\sigma_{p}^{2}}\right\}\right],$ (27)
@@
With $\sigma_{q}\leq\epsilon_{0}$, the convexity of two terms ($\operatorname{KL}$ and $l_{\mathcal{D}}$) holds. Hence we show by the Karush–Kuhn–Tucker theorem (Kjeldsen, 2000; Karush, 1939; Kuhn & Tucker, 1951) that the following two optimization problems are equivalent:
Minimization of Variational Free-Energy as the Lagrange-form optimization
$\min_{\sigma_{q}}\quad\operatorname{KL}[q(\operatorname{vec}({\bm{W}})|\sigma_{q})\parallel P(\operatorname{vec}({\bm{W}})|\sigma_{p})]+\tfrac{1}{\lambda}\cdot l_{\mathcal{D}}(\sigma_{q});$ (28)
The constrained-form optimization corresponding to Eqn. 28
@@
s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\epsilon.$
Since the $\operatorname{KL}$ term is monotonically decreasing when $\sigma_{q}\in[0,\sigma_{p})$, and due to the fact that $\sigma_{p}>\epsilon_{0}$, the optimization in Eqn. 29 is equivalent to:
Training-Free Bayesianization (TFB)
@@
s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\epsilon.$
Appendix B Implementation Details
B.1 Datasets

### v2

@@ 5.5 TFB Beyond the Naive LoRA
Table 4: Performance of different LoRA-like PEFTs on the combined dataset.
@@
Our proposed TFB is general, compatible with various LoRA-based methods that use different initialization strategies [41], parameter sharing schemes [35], and optimization approaches [67]. We evaluate TFB on two representative LoRA variants (see Appendix F.8 for details):
  • VeRA [35]: Uses shared low-rank matrices ${\bm{B}}$ and ${\bm{A}}$ across layers, with the layer-specific trainable scalar vector ${\bm{d}}$ and the bias vector ${\bm{b}}$.
  • PiSSA [41]: Employs an alternative initialization while maintaining LoRA’s training process.
Table 4 shows the results, demonstrating TFB’s broad applicability: it substantially reduces ECE across all LoRA-like PEFT methods (e.g., from 5.11 to 1.44 for VeRA). Importantly, it maintains strong ACC and NLL with minimal performance degradation, validating the effectiveness of low-rank isotropic Gaussian distributions for variational inference of LLMs.
5.6 Improving Inference-Time Efficiency of TFB
Table 5: Performance of Last-Layer TFB (LL TFB) applied to the combined dataset.
@@ 5.6 Improving Inference-Time Efficiency of TFB
Inspired by Harrison et al. [19], in this section, we study the compactibility of TFB and Last-Layer Bayesianization to speed up inference. Specifically, we employ the same standard deviation $\sigma_{q}^{*}$ as identified through the full-model TFB, but apply the Beyesianization to the last layer only. By limiting inference to sampling from the variational distribution of the last layer, the outputs of earlier layers can be reused, significantly improving inference speed. Although restricting Bayesianization to the last layer sacrifices some richness of the variational distribution family, this approach allows for a higher number of posterior samples, resulting in more accurate approximations.
Table 5 shows the results. When only $N=10$ samples are used, the uncertainty estimation of last-layer Bayesianization performs worse than that of full-model Bayesianization. Nevertheless, the improved posterior estimation with $N=100$ samples enables last-layer Bayesianization to achieve better performance. For a comprehensive analysis of its performance, please refer to Appendix F.9.
@@ 6 Conclusion
In this paper, we introduce Training-Free Bayesianization (TFB), a novel framework that transforms trained LoRA adapters into Bayesian ones without additional training. By systematically searching for the maximally acceptable variance in the weight posterior within a family of low-rank isotropic Gaussian distributions, TFB provides a practical solution to uncertainty estimation in LLMs. Our theoretical analysis shows that TFB’s variance maximization process is equivalent to generalized variational inference under mild conditions. Our empirical results verify its superior performance across various settings and model configurations. Our framework’s simplicity and effectiveness, requiring only an anchor dataset for search, makes it widely applicable across different domains. As LLMs continue to evolve, TFB represents a significant step toward more reliable and uncertainty-aware AI systems, paving the way for future research in adaptive and trustworthy machine learning.
@@
In Appendix A, we present the full algorithmic description of TFB. In Appendix B, we present the limitations of TFB. Next, in Appendix C, we present a more detailed introduction to recent advances of Bayesian Low-Rank Adaptation. In Appendix D, we provide detailed proofs for all theorems presented in the main paper. In Appendix E, we describe our experimental methodology. Finally, in Appendix F, we present additional empirical results, including:
  • a visual study demonstrating TFB functions as a general Bayesian Neural Network (Appendix F.1),
  • computational complexity analysis (Appendix F.2),
  • anchor dataset size analysis (Appendix F.3),
  • TFB with unlabeled test data as the anchor data (Appendix F.4),
  • TFB with non-NLL evaluation metrics (Appendix F.5),
  • TFB with other single-parameter variational distribution families (Appendix F.6),
  • TFB with other LLM backbones than Llama3.1-8B (Appendix F.7),
  • TFB with other PEFT methods than LoRA (Appendix F.8),
  • improving inference-time efficiency of TFB with last-layer Bayesianization (Appendix F.9), and
  • the full results on a widely used LLM architecture Llama2-7B (Appendix F.10).
Appendix A TFB: Algorithm
{algorithm}
[…]
$\{{\bm{B}},{\bm{A}}\}$: Low-Rank Component;
$l$: Model Evaluation Metric;
$\epsilon$: Performance Change Tolerance;
[…]
Figure 1: Overview of our Training-Free Bayesianization, i.e., TFB (right) as well as comparison with existing methods such as LoRA (middle) and BLoB (left).
Appendix B Broader Impact and Limitations
Broader Impact. This research advances methods for making large language models more trustworthy and reliable through improved uncertainty estimation. While we focus on language models, the fundamental principles of our framework can enhance uncertainty quantification across the broader machine learning field. This wider applicability creates opportunities for improving model reliability and safety across diverse applications. To the best of our knowledge, there are no ethical or other concerns that need to be addressed.
Limitations.TFB is subject to several limitations. First, our approach relies on the availability of an anchor dataset for determining search range and stopping criteria. Although this dataset doesn’t require supervision or prior use in LoRA training, its quality and representativeness could impact the effectiveness of uncertainty estimation. Second, by constraining the family of full-weight posteriors to low-rank isotropic Gaussian distributions, TFB may not capture more complex uncertainty patterns that could be present in the data. At first blush, this seems to imply a trade-off between computational efficiency and model expressiveness. However, in practice, the trade-off may not be necessary as TFB can often enjoy both computational efficiency and model expressiveness, getting the best of both worlds. Given TFB’s proven effectiveness and inherent simplicity, we recommend implementing TFB as the initial approach when developing reliable LLMs with existing LoRA adapters. If TFB fails to meet specific requirements, practitioners can then consider alternative expensive Bayesian methods. Finally, while we have demonstrated the effectiveness of TFB in various settings, its performance in more complex generation tasks requires further investigation. Future work could explore extending the framework to handle more sophisticated language generation scenarios and broader applications.
Appendix C Related Work
Bayesian Low-Rank Adaptation. The Bayesian framework provides a powerful approach for capturing and estimating uncertainty by defining prior distributions and approximating posterior distributions over the parameter space [46, 21, 14, 55, 18]. However, modeling parameter distributions across the entire parameter space during fine-tuning introduces significant computational overhead [13, 68]. To address this challenge, recent research has explored combining Bayesian methods with Parameter-Efficient Fine-Tuning (PEFT) techniques to improve the efficiency of uncertainty estimation. Several notable approaches have emerged in this direction. Wang et al. [60] and Balabanov & Linander [4] demonstrate improved performance by training multiple LoRA modules and ensemble their predictions during inference. Taking a different approach, Yang et al. [66] applies a Kronecker-factorized Laplace approximation to fine-tuned LoRA parameters. More recently, BLoB [61] advances the field by simultaneously estimating both the mean and covariance of LLM parameters within a single fine-tuning stage, leading to substantial performance improvements. Our proposed training-free Bayesianization represents a significant departure from these existing methods. Unlike approaches that require re-training [14, 60, 4, 61] or rely on continued training and gradient estimation [66], our method achieves uncertainty estimation without any additional training steps, substantially improving the simplicity and efficiency for Bayesian learning of LLMs.
Appendix D Proof of Theorems
Lemma D.1.
@@ Appendix D Proof of Theorems
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}].$
Theorem 4.1 (Equivalent Variational Distribution of the Full Weight ${\bm{W}}$ in TFB).
With the pre-trained weight matrix ${\bm{W}}_{0}\in\mathbb{R}^{m\times n}$, the low-rank weight update matrix $\{{\bm{B}}^{\prime}\in\mathbb{R}^{m\times r},{\bm{A}}^{\prime}\in\mathbb{R}^{r\times n}\}$ transformed from the given matrices $\{{\bm{B}},{\bm{A}}\}$ following Eqn. 4 and 5, suppose that the variational distribution of ${\bm{A}}^{\prime}$ is Gaussian $q({\bm{A}}^{\prime}|{\bm{\theta}})=\prod_{ij}{\mathcal{N}}(A_{ij}|M_{ij},\Omega_{ij}^{2})$, where ${\bm{M}}=[M_{ij}=A^{\prime}_{ij}]\in\mathbb{R}^{r\times n}$ is its mean and ${\bm{\Omega}}=[\Omega_{ij}]\in\mathbb{R}^{r\times n}$ is the standard deviation calculated as in Eqn. 7. The equivalent variational distribution $q(\operatorname{vec}({\bm{W}})|\sigma_{q})$ defined on the full weight matrix ${\bm{W}}$ is
${\bm{\Sigma}}_{q}$ $=\sigma_{q}^{2}{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}.$
Proof.
We have the following lemma from BLoB that calculates the covariance matrix of a given low-rank Bayesianization scheme $\{{\bm{B}},{\bm{A}},{\bm{\Omega}}\}$ [61].
Based on the assumption outlined in Eqn. 4, 5, and 7, we have the following properties about ${\bm{B}}^{\prime}$, ${\bm{M}}$, and ${\bm{\Omega}}$ of TFB:
@@
where ${\bm{U}}^{\top}{\bm{U}}={\bm{I}}_{r},{\bm{U}}{\bm{U}}^{\top}=\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}.$ (18)
@@
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}^{\prime}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})][{\bm{I}}_{n}\otimes{\bm{B}}^{\prime\top}]$ (19)
$=[{\bm{I}}_{n}\otimes{\bm{B}}^{\prime}][{\bm{I}}_{n}\otimes\operatorname{diag}(1/{\bm{d}})^{2}][{\bm{I}}_{n}\otimes{\bm{B}}^{\prime\top}]$ (20)
@@
$=\sigma_{q}^{2}{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix},$ (23)
@@
Proposition D.1.1.
The function $\operatorname{proj}(\cdot)$ defined in Eqn. 3 projects the full-dimensional isotropic Gaussian to the low-rank subspace of LoRA. It can be formulated as
$\operatorname{proj}(\sigma_{q}^{2}{\bm{I}}_{mn})$ $={\bm{P}}(\sigma_{q}^{2}{\bm{I}}_{mn}),$ (24)
$\text{where}\quad{\bm{P}}$ $={\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}.$ (25)
Proof.
By Theorem 4.1, we have
${\bm{P}}{\bm{I}}_{mn}$ $={\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}.$ (26)
Hence it is trivial to have ${\bm{P}}={\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}.$
Theorem 4.2 (TFB as Generalized Variational Inference).
Suppose the evaluation metric $l_{\mathcal{D}}(\sigma_{q})$ defined following Assumption 4.1 is locally convex within the range of $\sigma_{q}\in[0,\epsilon_{0})$. Suppose the approximate distribution of ${\bm{W}}$ given $\sigma_{q}$ is defined following Theorem 4.1. Suppose we have the prior distribution $P(\operatorname{vec}({\bm{W}}))={\mathcal{N}}(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{p},{\bm{\Sigma}}_{p})$, where ${\bm{\mu}}_{p}={\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}})$, and ${\bm{\Sigma}}_{p}=\sigma_{p}^{2}{\bm{I}}$ with $\sigma_{p}>\epsilon_{0}$. Then for $\forall\lambda>0$, $\exists\widetilde{\epsilon}$, s.t. the following two optimization problems (i) Generalized Variational Inference [7, 22, 30]
@@
s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\widetilde{\epsilon},$
are equivalent, i.e., the two optimization problems have the same optimal solution, where $\lambda$ is the regularization coefficient of the KL-divergence.
@@
First we prove the KL divergence term is convex w.r.t. $\sigma_{q}$. For two Gaussian distributions $q$ and $p$ whose covariance matrices ${\bm{\Sigma}}_{q}\in\mathbb{R}^{d\times d}$ and ${\bm{\Sigma}}_{p}\in\mathbb{R}^{d\times d}$ are both full-rank, with their means as ${\bm{\mu}}_{q}\in\mathbb{R}^{d}$ and ${\bm{\mu}}_{p}\in\mathbb{R}^{d}$, we have their KL-divergence as
@@
$q(\operatorname{vec}({\bm{W}})|\sigma_{q})$ $={\mathcal{N}}\left(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}}),{\bm{\Sigma}}_{q}=\sigma_{q}^{2}{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}\right),$ (30)
$P(\operatorname{vec}({\bm{W}})|\sigma_{p})$ $={\mathcal{N}}\left(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{p}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}}),{\bm{\Sigma}}_{p}=\sigma_{p}^{2}{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}\right).$
Substituting Eqn. 30 back into Eqn. 29, we have
$\operatorname{KL}[q(\operatorname{vec}({\bm{W}})|\sigma_{q})\|P(\operatorname{vec}({\bm{W}}|\sigma_{p}))]$ $=\tfrac{nr}{2}\left[\log(\sigma_{p}^{2})-1+\left\{-\log(\sigma_{q}^{2})+\tfrac{\sigma_{q}^{2}}{\sigma_{p}^{2}}\right\}\right],$ (31)
@@
With $\sigma_{q}\leq\epsilon_{0}$, the convexity of two terms ($\operatorname{KL}$ and $l_{\mathcal{D}}$) holds. Hence we show by the Karush–Kuhn–Tucker theorem [32, 29, 37] that, for any given $\lambda$ there exists $\widetilde{\epsilon}$ such that the following two optimization problems are equivalent:
  1. 1.
    Minimization of generalized variational inference in the Lagrange-form optimization
    $\min_{\sigma_{q}}\quad\operatorname{KL}[q(\operatorname{vec}({\bm{W}})|\sigma_{q})\parallel P(\operatorname{vec}({\bm{W}})|\sigma_{p})]+\tfrac{1}{\lambda}l_{\mathcal{D}}(\sigma_{q});$ (32)
  2. 2.
    The constrained-form optimization corresponding to Eqn. 32
@@
    s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\widetilde{\epsilon}.$
Since the $\operatorname{KL}$ term is monotonically decreasing when $\sigma_{q}\in[0,\sigma_{p})$, and due to the fact that $\sigma_{p}>\epsilon_{0}$, the optimization in Eqn. 33 is equivalent to our final Training-Free Bayesianization (TFB):
@@
s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\widetilde{\epsilon}.$
Appendix E Implementation Details
E.1 Datasets